A certain oscillator has equation of motion
x¨ + 4x = 0.
The particle is initially at the point x =

3 when it is projected towards the origin with a speed 2. Determine the displacement, x(t), of the particle
1
Expert's answer
2020-11-17T11:25:13-0500

Answer

Given equation

"\frac{d^2x}{dt^2}=-4x"

Given equation is of the simple harmonic motion so

"\omega=\sqrt4=2rad/s"

Displacement in shm is given by

"x(t) =x_ocos(wt+\phi)"


At t=0sec

"\sqrt{3}=Acos(2\times0+\phi) \\\sqrt{3}=Acos\phi" . . . ... Eq. 1

And velocity v=2m/s

So velocity can be written as for shm

"v=-x_o\omega sin(\omega t+\phi)"

At t=0sec

"-2=-2x_osin(2\times0+\phi) \\x_osin\phi=1" .. Eq. 2

By solving equation 1 and 2

"tan\phi=\frac{1}{\sqrt{3}}"

"\phi=\frac{\pi}{6}"


And "x_o=2"

So equation can be Written as

"x(t) =2cos(2t+\frac{\pi}{6})"